For the given Parabola (y-5)=16(x+2)^2 , determine: a. vertex b.P c. focus d. equation of directrix

  1. For the given Parabola (y-5)=16(x+2)^2 , determine:

  2. a. vertex

  3. b.P

  4. c. focus

  5. d. equation of directrix
Answer #1

2- ५-5.- 10612) • ४ = 16(1- (9)+5 it is in the firm of y = 4 (4-) +K where, (A) vaten (hik) = (-2,S) (19) = ( 0, 6) 1303; 16